# Sparse Covariance Neural Networks

### v1

@@ 4.3 Stochastic sparsification strategies
Absolute covariance values (ACV). Despite being dense, the covariance matrix may contain some values of lower magnitude that may be reminiscent of spurious correlations. We want therefore to drop these small covariance values with a high probability to save computational cost while keeping large covariance values to maintain the useful dependencies. Thus, we define the probability as $p_{ij}=|\hat{c}_{ij}|/\hat{c}_{\text{max}}$, where $\hat{c}_{\text{max}}=\max_{i,j}|\hat{c}_{ij}|$. Fig. 2 (left) shows that the stability term $q_{in}$ within $Q$ [cf. Thm. 4] is small when the covariance value $\hat{c}_{in}$ approaches zero or the maximal value $\hat{c}_{\max}$. This is because small covariance values have little impact on the VNN stability even if they are more likely to be dropped, and large covariance values (close to $\hat{c}_{\text{max}}$) are less likely to be dropped though their removal would affect stability more. It peaks around $\hat{c}_{in}=2/3\hat{c}_{\text{max}}$, which identifies an interval in which ACV may have a notable impact on the stability.
@@
Ranked covariance values (RCV). While ACV improves efficiency by preserving stronger correlations, it does not allow any control over the amount of sparsification (e.g., if all covariance values are high, very few are dropped). To overcome this, we define a set of probability values with a desired mean $p$ and assign them to covariance values based on their positions in this absolute ranking. Formally, consider an ordered set of probabilities $\mathcal{P}=\{p^{\prime}_{1},\dots,p^{\prime}_{N^{\prime}}\}$, where $p^{\prime}_{i}\leq p^{\prime}_{i+1}$, and each $p^{\prime}_{i}$ is sampled from $\mathcal{N}(p,\sigma)$ with $\sigma=\min((1-p)/3,p/3))$ (such that the number of values not in $[0,1]$, which we clip to the interval, is negligible) and $N^{\prime}$ is the number of probability values for assignment. We set $p_{ij}=p^{\prime}_{k}$ where $k=|\{c_{lm}:|c_{lm}|<|c_{ij}|;l,m=1,\dots,N\}|$ is the position of $c_{ij}$ in the ordered ranking of absolute covariances. Consequently, the expected percentage of dropped covariances is $1-p$, which allows controlling the sparsification level at the risk of removing useful covariances. In practice, the value of $p$ can be treated as a hyperparameter and tuned through cross-validation or on a validation set based on the desired accuracy and stability.
@@
Fig. 2 (right) shows how the term $Q$ changes for ACV and RCV with different means of covariance value distribution. RCV with a smaller $p$ corresponds to lower stability especially when the covariance values are high, because they are dropped regardless of their value. ACV, instead, maintains a consistent level of stability since it balances the covariance magnitudes and the dropping probabilities.
@@
Computational complexity. The expected computational complexity of a VNN layer with stochastic sparsification is of the order $\mathcal{O}(\mathbb{E}[\|\mathbf{\tilde{C}}\|_{0}]KF_{\text{in}}F_{\text{out}})$, where $\mathbb{E}[\|\mathbf{\tilde{C}}\|_{0}]$ is the expected number of non-zero elements of the sparsified covariance. For ACV, we have that $\mathbb{E}[\|\mathbf{\tilde{C}}\|_{0}]=\hat{c}_{\text{mean}}(\|\mathbf{\hat{C}}\|_{0}-N)+N$, where $\hat{c}_{\text{mean}}=1/N^{2}\sum_{i,j=1,\dots,N,i\neq j}|\hat{c}_{ij}|$, so the complexity depends on the data. Instead, for the RCV we have $\mathbb{E}[\|\mathbf{\tilde{C}}\|_{0}]=p(\|\mathbf{\hat{C}}\|_{0}-N)+N$, so the complexity can be controlled through $p$.
@@
The stochastic sparsification with RCV generalizes the hard thresholding approach in Sec. 3, which can be recovered by setting probabilities as $p_{ij}=1$ if $|c_{ij}|>\tau/\sqrt{t}$, 00 otherwise.
@@
Both thresholding and stochastic sparsification provide symmetric covariance estimates but do not necessarily preserve positive semidefiniteness (PSD), which is common in related works (Bickel & Levina, 2008b; Deshpande & Montanari, 2016; Li et al., 2021; Liao et al., 2022) and places our techniques within the scope of the literature. Bickel & Levina (2008b) also provide a sufficient condition for PSD under hard-thresholding, which we report in Appendix G.
Figure 2: Insights into stability for stochastic sparsification. (Left) Stability bound elements $q_{in}=\hat{c}_{in}^{2}(1-|\hat{c}_{in}|/\hat{c}_{\text{max}})$ with $Q=\sum_{i=1}^{N}\sum_{n=1}^{N}q_{in}$ for ACV [cf. Thm. 4]. (Right) Stability term $Q$ [cf. Thm. 4] for covariance values distributions with different means and stochastic sparsifications.
5 Numerical Results
@@
We corroborate the proposed methods with experiments on real and synthetic datasets, targeting the following objectives: validate our theoretical results on (O1) thresholding when the true covariance matrix is sparse and (O2) stochastic sparsification for a generic true covariance; (O3) compare S-VNNs and VNNs on real datasets. We provide details and additional results in Appendix I.
5.1 Stability of sparse VNN
@@
Experimental setup – (O1)-(O2). We generate synthetic data with a controlled covariance by sampling data points $\mathbf{x}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{C})$. Then, we create regression targets as $y_{i}=\mathbf{w}^{\mathsf{T}}\mathbf{x}_{i}+u$, where $\mathbf{w}$ is a vector with elements $w_{j}\sim\text{Uniform}(0,1)$ and $u\sim\mathcal{N}(0,3)$ is a noise term. We generate 1000 samples of size $N=100$ and divide the data into train/validation/test splits of size 80%/10%/10%.
@@
Sparse true covariance – (O1). We evaluate the stability of S-VNN on a synthetic dataset with sparse covariance (SparseCov). We train a VNN with the true covariance and test it with different covariance estimates obtained with a varying number of samples. We compare the original sample covariance, hard thresholding (Hthr), and soft thresholding (Sthr). We use PCA-SVM as baseline, i.e., we transform data with PCA followed by an RBF-SVM for regression. Fig. 3 (left) reports the MAE w.r.t. the number of samples for covariance estimation. Overall, hard and soft thresholding provide better covariance estimates, ultimately, improving the performance (i.e., lower MAE) and increasing the stability compared to the nominal VNN. VNNs are more stable than PCA-based models, because their performance is less affected by covariance perturbations as analyzed in (Sihag et al., 2022). These results support our theoretical discussion in Sec. 3, i.e., thresholding allows VNNs to maintain stability in sparse covariance settings.
@@
Generic true covariance – (O2). We assess the stochastic sparsification using synthetic datasets with large covariance values (LargeCov) and with small covariance values (the synthetic linear regression dataset in (Sihag et al., 2022, Appendix E.4) with tail=0.2, here called SmallCov); see Sec. I.1 for the distributions of the covariance values. We train a VNN with the true covariance matrix and test it with the stochastically sparsified versions. Fig. 3 (center) and (right) compare the ACV and RCV with different values of $p$. For LargeCov, we see that a high sparsification leads to greater instability (RCV for small $p$) as large covariance values are dropped which hinders performance. For SmallCov, instead, all sparsification approaches affect stability only lightly as removing small covariance values does not lead to performance changes. ACV remains stable because it sparsifies little even when the covariance values are large. In turn, this corroborates our theoretical observations in Sec. 4.1.
Figure 3: Stability of VNN. (Left) MAE with hard and soft thresholding on SparseCov (standard deviations are of order at most $10^{-2}$ for VNNs and $10^{-1}$ for PCA). (Center) MAE for stochastic sparsification on LargeCov. (Right) MAE for stochastic sparsification on SmallCov.
@@
Figure 4: Accuracy and time for a forward pass for sparse and dense VNNs. From left to right, the results are for Epilepsy, CNI, MHEALTH and Realdisp. RCV results are for $p=0.75,0.5,0.25$ from top to bottom, and for Hthr and Sthr we report results for 3 different thresholds to achieve analogous sparsification. Legend is shared.
5.2 Sparse VNN on real data
Experimental setup – (O3). We consider four real datasets of brain measurements and human action recognition data. Epilepsy (Kramer et al., 2008) consists of an electrocardiography time series equally split before and after an epileptic seizure event. We perform binary classification to predict whether a time sample was recorded before or after the seizure. CNI (Schirmer et al., 2021) contains resting-state fMRI time series of patients with attention deficit hyperactivity disorder (ADHD) and neurotypical controls (NC), which we use as labels for binary classification. MHEALTH (Banos et al., 2014a) and Realdisp (Banos et al., 2014b) contain measurements of wearable devices placed on subjects performing different actions. The goal is to classify the action performed by each subject. Additional details are reported in Sec. I.1. We compare various sparsification approaches (thresholding, RCV, ACV) against the dense VNN. For completeness, we report additional results for various PCA-based classifiers and GraphSAGE in Sec. I.3, although our objective is not to achieve state-of-the-art performance on these datasets but to reduce the impact of spurious correlations while improving computational efficiency of VNNs. For Hthr and Sthr, we set the threshold to achieve a level of sparsification comparable to RCV with $p=0.75,0.5,0.25$ to ease the comparison.
Results – (O3). From Fig. 4, we see that the S-VNNs improve substantially on the computation time and increase the accuracy w.r.t. the nominal VNN. This is particularly emphasized for Epilepsy, CNI, and MHEALTH, where most S-VNNs perform better than dense VNN, likely because the sample covariance matrices in these datasets contain spurious correlations. On Realdisp, the accuracy improvement is smaller, indicating that in this dataset most correlations carry useful information to solve the task. Both for thresholding and RCV, changing the value of the threshold or $p$ does not lead to large changes in accuracy, meaning that the majority of covariance values are not relevant for performance, while it affects the time efficiency of the model by increasing sparsity. Finally, on 3 out of 4 datasets, stochastic sparsification achieves a slightly superior performance than deterministic thresholding, likely due to the regularization effect of different realizations of a sparsified covariance over different runs in a form akin to dropout-like techniques (Papp et al., 2021; Rong et al., 2020).
6 Related Works

### v2

@@ IV-B Sparsification Probability
Absolute covariance values (ACV). Due to spurious correlations, the sample covariance matrix may contain some values of lower magnitude. We want to drop these small values with a high probability to save computational cost while keeping large covariance values to maintain the useful data correlations. Thus, we define the probability as $p_{ij}=|\hat{c}_{ij}|/\hat{c}_{\text{max}}$, where $\hat{c}_{\text{max}}=\max_{i,j}|\hat{c}_{ij}|$. Fig.˜4 (left) shows that the stability term $q_{in}=\hat{c}_{in}^{2}(1-|\hat{c}_{in}|/\hat{c}_{\text{max}})$ with $Q=\sum_{i=1}^{N}\sum_{n=1}^{N}q_{in}$ [cf. Thm.˜5] is small when the covariance value $\hat{c}_{in}$ approaches zero or the maximal value $\hat{c}_{\max}$. This is because small covariance values have little impact on the S-VNN stability even if they are more likely to be dropped, and large covariance values (close to $\hat{c}_{\text{max}}$) are less likely to be dropped though their removal would affect stability more.
@@
Ranked covariance values (RCV). While ACV improves efficiency by preserving stronger correlations, it does not allow to control the amount of sparsification (e.g., if all covariance values are high, very few are dropped). To overcome this, we define a set of probabilities with a desired mean $p$ and assign them to covariance values based on their positions in the absolute ranking. Formally, consider an ordered set of probabilities $\mathcal{P}=\{p^{\prime}_{1},\dots,p^{\prime}_{N^{\prime}}\}$, where $p^{\prime}_{i}\leq p^{\prime}_{i+1}$, and each $p^{\prime}_{i}$ is sampled from $\mathcal{N}(p,\sigma)$ with $\sigma=\min((1-p)/3,p/3))$ (such that the number of values not in $[0,1]$, which we clip to the interval, is negligible) and $N^{\prime}$ is the number of probability values for assignment. We set $p_{ij}=p^{\prime}_{k}$ where $k=|\{c_{lm}:|c_{lm}|<|c_{ij}|;l,m=1,\dots,N\}|$ is the position of $c_{ij}$ in the ordered ranking of absolute covariances. Consequently, the expected percentage of dropped covariances is 1p1-p, which allows controlling the sparsification level at the risk of removing useful covariances. The value of $p$ can be treated as a hyperparameter and tuned through cross-validation or on a validation set based on the desired accuracy and stability.
@@
Fig.˜4 (right) shows how the term $Q$ changes for ACV and RCV with different means of covariance value distribution. RCV with a smaller $p$ corresponds to a lower stability especially when the covariance values are high, because they are dropped regardless of their value. ACV, instead, maintains a consistent level of stability since it balances the covariance magnitudes and the dropping probabilities, but it does not allow to control the level of sparsification.
@@
Computational complexity. The expected computational complexity of a S-VNN layer with stochastic sparsification is of the order $\mathcal{O}(\mathbb{E}[\|\mathbf{\tilde{C}}\|_{0}]KF_{\text{in}}F_{\text{out}})$, where $\mathbb{E}[\|\mathbf{\tilde{C}}\|_{0}]$ is the expected number of non-zero elements of the sparsified covariance. For ACV, we have that $\mathbb{E}[\|\mathbf{\tilde{C}}\|_{0}]=\hat{c}_{\text{mean}}(\|\mathbf{\hat{C}}\|_{0}-N)+N$, where $\hat{c}_{\text{mean}}=1/(N^{2}-N)\sum_{i,j=1,\dots,N,i\neq j}|\hat{c}_{ij}|$, so the computation complexity depends on the data. For RCV, we have $\mathbb{E}[\|\mathbf{\tilde{C}}\|_{0}]=p(\|\mathbf{\hat{C}}\|_{0}-N)+N$, so the computation complexity can be controlled through the desired mean $p$.
@@
The stochastic sparsification can be considered as an extension of the hard thresholding in Sec. III, which allows to control the degree of sparsification through the sampling probabilities $\{p_{ij}\}$. It reduces to the hard thresholding by setting probabilities $p_{ij}=1$ if $|c_{ij}|>\tau/\sqrt{t}$, and 0 otherwise.
@@
Both hard/soft thresholding and stochastic sparsification provide symmetric covariance estimates, but do not necessarily preserve positive semidefiniteness (PSD). While this is common [25, 18, 37, 5], if PSD estimates are of interest one can perform diagonal loading $\mathbf{\hat{C}}+\delta\mathbf{I}$ for a $\delta\geq 0$ that ensures all eigenvalues are non-negative. Moreover, [25] provides a sufficient condition for PSD under hard-thresholding. Specifically, consider a hard-thresholded sample covariance matrix $\mathbf{\bar{C}}$ as per Def.˜2 and the sample covariance matrix $\mathbf{\hat{C}}$. The sparsified $\mathbf{\bar{C}}$ is PSD if $\|\mathbf{\bar{C}}-\mathbf{\hat{C}}\|\leq\epsilon$ and $\lambda_{\textnormal{min}}(\mathbf{\hat{C}})>\epsilon$ for an $\epsilon>0$, where $\lambda_{\textnormal{min}}(\cdot)$ computes the smallest eigenvalue. That is, if the true covariance matrix is sparse, then the term $\|\mathbf{\bar{C}}-\mathbf{\hat{C}}\|$ decreases as the number of samples $t$ increases and the sparse estimate is more likely to be PSD.
V Numerical Results
@@
We evaluate the proposed S-VNN with different sparsification strategies through experiments on both real and synthetic datasets, targeting the following objectives: (O1) validate the sparsification strategies and theoretical results when the true covariance is sparse; (O2) validate the sparsification strategies and theoretical results when the true covariance matrix is dense; (O3) evaluate the performance and computation efficiency of the S-VNN over several baselines on real-world data.
V-A Datasets
We consider one controlled scenario and four real datasets.
@@ V-A Datasets
Synthetic scenario. We evaluate S-VNNs on a synthetic regression task, analogously to [11]. We generate synthetic data with a controlled covariance by sampling data points $\mathbf{x}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{C})$. Then, we create regression targets as $y_{i}=\mathbf{w}^{\mathsf{T}}\mathbf{x}_{i}+u$, where $\mathbf{w}$ is a vector with elements $w_{j}\sim\text{Uniform}(0,1)$ and $u\sim\mathcal{N}(0,3)$ is a noise term. We generate 1000 samples of size $N=100$ and divide the data into train/validation/test splits of size 80%/10%/10%. We consider three different conditions: we use a sparse $\mathbf{C}$ (SparseCov), a dense $\mathbf{C}$ with large values (LargeCov) and, for the case of a dense $\mathbf{C}$ with small values (SmallCov), we use the synthetic linear regression dataset in [11, Appendix E.4] with tail=0.2. We show the covariance distributions of LargeCov and SmallCov in Fig.˜5.
TABLE I: Details on datasets.
Dataset Nodes Node Classes Samples
features (train/valid/test)
MHEALTH 23 128 12 5357 (3239/1028/1090)
Realdisp 117 128 10 5660 (2683/1258/1719)
ADNI1 68 1 NA 801 (481/160/160)
ADNI2 68 1 NA 1142 (685/228/229)
Figure 5: Normalized absolute covariance values distributions for real and synthetic datasets.
Real datasets. We also consider four real-world datasets of brain data and human action recordings for S-VNN stability and performance evaluation. Table˜I summarizes the main information about the datasets and Fig.˜5 shows their empirical normalized covariance values distribution.
  • MHEALTH [38] contains measurements of wearable devices placed in the chest, right wrist and left ankle of 10 subjects performing 12 different actions. Following [39], we use recordings from different subjects for train, validation and test. Specifically, we use subjects 6, 10 for validation, 2, 9 for test and the rest for training. We do not use the ECG measurements placed in the chest. The goal is to classify the action performed by the subject. The sampling rate is 50 Hz.
  • Realdisp [40] contains data of 17 subjects performing 33 actions. We use subjects 10, 11, 12 for validation; subjects 13, 14, 15, 16, 17 for test and the remaining for training. We use the 10 most common activity labels: walking, jogging, running, cycling, elliptic bike, trunk twist (arms outstretched), rowing, knees (alternatively) bend forward, waist bends forward, trunk twist (elbows bended). We segment the data in MHEALTH and Realdisp creating sliding windows of size 128 (2.56 seconds) for measurements relative to the same activity with 50% overlap.
  • ADNI1 and ADNI2 [41] correspond to two different phases of the ADNI project containing cortical thickness measures of patients affected by Alzheimer’s disease and healthy control patients. The phases differ in the collection protocols and historical periods. We download the cortical thickness measures for both datasets (i.e., the outputs of the FreeSurfer [42] processing of the brain MRI recordings) from https://ida.loni.usc.edu/. For each patient, we consider the first scan available, generally corresponding to the screening visit. The task is a regression problem that aims to classify the chronological age of the patient at the time of the scan from its cortical thickness measures. This is a task of high interest to identify neurodegenerative diseases from early brain aging, since cortical thickness generally decreases with age. Correlations among brain regions play a relevant role in characterizing cortical thickness patterns [13, 43, 44, 45], making this task relevant for S-VNNs.
Figure 6: Stability of hard and soft thresholded S-VNNs, nominal VNN and PCA-SVM with dense and sparse covariance on regression task for a synthetic dataset with sparse true covariance. (Above) Regression performance in terms of Mean Absolute Error. (Below) Embedding difference (i.e., $\|\mathsf{\Phi}(\mathbf{x},\mathbf{\bar{C}},\mathcal{H})-\mathsf{\Phi}(\mathbf{x},\mathbf{C},\mathcal{H})\|$) between S-VNN/VNN/PCA with true and estimated covariance. Standard deviations in our results are of order at most $10^{-2}$ for VNNs and $10^{-1}$ for PCA.
V-B Stability of S-VNN
@@
Sparse true covariance on synthetic data – (O1). We evaluate the proposed hard thresholding (Hthr) and soft thresholding (Sthr) using the SparseCov dataset. We train a VNN with the true covariance and test it with different covariance estimates under varying number of samples. We consider the nominal VNN with the original sample covariance and the PCA-SVM as baselines for comparison. For PCA-SVM, we transform data with PCA and we then apply an RBF-SVM for regression. Fig.˜6 reports the Mean Absolute Error (MAE) on the downstream task and the embedding difference between the models operating on the true and sample covariances, w.r.t. the number of samples for covariance estimation. Overall, hard and soft thresholding provide better covariance estimates, ultimately, improving the performance (i.e., lower MAE) and increasing the stability (i.e., lower embedding difference) compared to the nominal VNN. S-VNN and VNN are more stable than PCA-SVM, because their performance is less affected by covariance perturbations as analyzed in Sec. II-B. These results support our theoretical discussion in Sec.˜III, i.e., thresholding allows VNNs to maintain or even improve stability in sparse settings.
@@
Generic true covariance on synthetic data – (O2). We assess the stochastic sparsification using LargeCov and SmallCov. We train a VNN with the true covariance matrix and test it with stochastically sparsified sample covariance matrices. Fig.˜7 shows the performance of the models operating over the true covariance and the sample covariance with the proposed ACV and RCV of different sparsification means $p$. For LargeCov, we see that a high sparsification leads to greater degradation (RCV for small $p$) as large covariance values are dropped which hinders performance. For SmallCov, instead, all sparsification approaches affect MAE only lightly as removing small covariance values does not lead to significant performance changes. ACV maintains stable performance for both LargeCov and SmallCov because it sparsifies little when the covariance values are large, allowing, however, for relatively small computational improvements. RCV, instead, allows for consistent sparsification levels and faster computation at the cost of more degradation when covariance values are large. These results corroborate our stability analysis in Theorem 5 and theoretical observations in Sec.˜IV-A.
Stochastic sparsification on real datasets – (O2). Fig.˜8 shows the stability of S-VNNs under stochastic sparsification with different strategies on the 4 real-world datasets. On Realdisp, which contains mostly small covariance values, the effect of sparsification is minimal, as the accuracy of S-VNNs remains consistent for most values of $p$. This indicates that most of the small correlations present in this dataset are not relevant for the downstream task. On ADNI1 and ADNI2, the covariance sparsification leads to a significant impact on the regression error for $p<0.5$, as these datasets contain large covariance values (cf. Fig. 5) whose removal affects the downstream performance. On MHEALTH, instead, the accuracy of S-VNN drops significantly for $p<0.5$. This indicates that, despite the covariance values being small, they still carry relevant information for the downstream task. Moreover, the classification task for MHEALTH contains 12 classes, making it more challenging than the age regression task on brain datasets, which may lead to lower model stability.
Figure 7: Regression performance in terms of MAE for S-VNNs with different stochastic sparsification techniques. (Left) LargeCov and (Right) SmallCov.
Figure 8: Stability of S-VNNs (in terms of average accuracy and standard deviation) on MHEALTH, Realdisp, ADNI1 and ADNI2 (from left to right) with different stochastic covariance sparsification techniques.
@@
Figure 9: MAE/Accuracy (%) and time for a forward pass on the test set for sparse and dense VNNs. From left to right, the results are for MHEALTH Realdisp, ADNI1 and ADNI2. RCV results are for $p=0.75,0.5,0.25$ from top to bottom, and for Hthr and Sthr we report results for 3 different thresholds to achieve analogous sparsification.
V-C Performance of S-VNN
Experimental setup – (O3). We compare S-VNNs with different sparsification strategies (i.e., hard thresholding, soft thresholding, RCV, ACV) against four categories of baselines: (i) the nominal VNN [11] which does not enforce sparsity; (ii) an MLP on the concatenation of raw node features that does not account for covariance information; (iii) an MLP on the concatenation of node features preprocessed by PCA, sparse PCA [46] and kernel PCA [47], which consider covariance information via PCA preprocessing; (iv) a GNN operating on a sparse graph obtained via neighbor sampling [48], where sparsification is not aware of covariance values but random. Our objective is to highlight the impact of sparsification, which reduces spurious correlations and improves computational efficiency of VNNs. For Hthr and Sthr, we compare 3 different thresholds that achieve a level of sparsification comparable to RCV with $p=0.75,0.5,0.25$ to compare their performance for analogous computational gains.
Results – (O3). From Fig.˜9, we see that the S-VNNs decrease substantially the computation time and improve the downstream performance compared with the VNN. The time improvements are evident on ADNI2 and especially the larger MHEALTH and Realdisp for most S-VNNs. On ADNI1, given its small size, the computation time is overall low for all models (if compared, for example, with ADNI2), and the impact of sparsification is less significant. Moreover, the larger computational time of Sthr compared to other models is due to its best hyperparameter configuration, which contains S-VNNs with 2 or 3 layers and order $K=3$, as opposed to the 1 layer and $K=1$ of the best configurations of other models. The Sthr with the same parameters of VNN has a computation time of $0.68\pm 0.03$ (faster than VNN).
In terms of downstream performance, the best MAE and accuracy are always obtained by S-VNNs, indicating that dense covariances contain spurious correlations that are detrimental for performance. This is particularly evident on ADNI2, where several sparse variants outperform VNN. For S-VNNs with both thresholding and RCV, changing the value of the threshold or the desired mean $p$ does not lead to large changes in accuracy, meaning that the majority of covariance values are not relevant for performance, while it affects the time efficiency of the model by increasing sparsity. This indicates that relatively higher sparsification, i.e., a larger threshold or a smaller $p$, may be a good choice in these cases.
Comparison with MLP, PCA and neighbor sampling – (O3).
Table˜II reports the results for S-VNNs, VNN (as in Fig.˜9) and additional baselines. On all datasets, the best downstream performance is attained by S-VNNs, indicating the importance of leveraging covariance information in a sparse manner. The MLP on the raw node features performs significantly worse than most other models on ADNI1, ADNI2 and Realdisp, highlighting the importance and effectiveness of considering covariance information as inductive bias. It is, however, faster than the other models as it does not involve any feature preprocessing. Preprocessing the data with PCA and variants generally improves performance, as the relevant information is effectively summarized, but it comes at the cost of higher computational time due to the expensive covariance eigendecomposition for standard PCA and even more complex optimization problems for sparse and kernel PCA. On MHEALTH and Realdisp, kernel PCA causes an out-of-memory error due to its dependence on the number of samples which is large for these datasets. Finally, the neighborhood sampling approach for sparsification leads to worse downstream performance than nominal VNNs, indicating that sparsification techniques that do not consider covariance values might lose relevant information.
Remark 3.
V-D Experimental details
Hyperparameters. On the synthetic datasets, we use VNNs with 2 layers, feature size 32 for LargeCov and SmallCov and 13 for SparseCov, we train for 50 epochs using Adam optimizer with learning rate 0.015 and weight decay $10^{-5}$. On real datasets, we perform a grid search for all models (VNN, S-VNNs and baselines) on all datasets among the following values: for VNN and S-VNNs, embedding size $F=32$, number of layers $L\in\{1,2,3\}$, dropout rate $dr\in\{0.0,0.3\}$, VNN order $K\in\{1,2,3\}$. We use the same parameter grid for the GCN except for $K$. We average the node embeddings and feed them to a 2-layer MLP as readout for the downstream task. We use batch normalization, Adam optimizer with learning rate 0.01 and weight decay $10^{-5}$ and we train for 1000 epochs with a patience of 100. For the MLP baseline (standalone or after PCA preprocessing), we search among three configurations: $L=2$ layers of size $32$, $L=2$ layers of size $256,128$, $L=3$ layers of size $256,128,128$. For the various PCA, we select the number of components $k\in\{10,20,50\}$. Table˜III reports the best hyperparameters after the search (for S-VNN, we report the configuration of the best-performing variant).
Implementation details. We repeat all experiments 5 times and, for stochastic experiments, we sample 10 different sparsified covariance matrixes. We report average results and standard deviations. Experiments are run on a AMD EPYC 7452 32-Core Processor with 10 GB of RAM, on a 13th Gen Intel Core i7-1365U with 16 GB of RAM and on an NVIDIA A40 GPU. We implement models using Pytorch and for neighborhood sampling we use Pytorch Geometric [49]. The sample covariance matrix $\mathbf{\hat{C}}$ is computed once over the complete training set and fixed during training, i.e., it is not affected by batching. For stability experiments, we use different sparsified covariances during test, whereas for the other experiments we keep the same matrix estimated on the training set.
TABLE II: Accuracy (%) and time (sec) for a forward pass on the test set on real datasets. Time results for ADNI1 and ADNI2 are $\cdot 10^{-3}$.
Accuracy (%) $\uparrow$ MAE $\downarrow$ Time (sec) $\downarrow$
MHEALTH Realdisp ADNI1 ADNI2 MHEALTH Realdisp ADNI1 ADNI2
[…]
TABLE III: Selected parameters on real datasets.
MHEALTH Realdisp ADNI1 ADNI2
VNN $L$ 3 2 1 3
drdr 0.3 0.3 0.0 0.0
$K$ 3 2 1 2
S-VNN $L$ 3 2 1 2
drdr 0.3 0.3 0.0 0.0
$K$ 3 3 1 2
NeighSampl $L$ 1 2 2 3
drdr 0.3 0.3 0.0 0.0
MLP $L$ 3 2 3 2
PCA+MLP $L$ 2 2 3 3
$k$ 23 50 10 10
SPCA+MLP $L$ 2 2 3 3
$k$ 20 20 10 10
KPCA+MLP $L$ NA NA 3 3
$k$ NA NA 10 10
VI Related Works
Our contribution relates to existing literature on sparse covariance estimation for PCA, stability analysis of graph and covariance neural networks, and on the sparsification strategies used for GNNs in general. In the sequel, we review these three streams and position our contribution accordingly.